OpenStudy (anonymous):

A triangular trellis has angles R, S, and T that measure 73°, 73°, and 34°, respectively. If the length of ST = 4y + 6 and the length of TR = 7y - 21, what is the value of y?

OpenStudy (anonymous):

picture?

OpenStudy (anonymous):

Theyr was no picture

OpenStudy (anonymous):

use law of sines

OpenStudy (anonymous):

Sine rule: \[\frac{ST}{\sin(R)} = \frac{TR}{\sin(S)}\] \[\frac{4y+6}{\sin(73^o)} = \frac{7Y-21}{\sin(73^o)}\] The rest should come easily.

OpenStudy (anonymous):

length ST=r length RS=t length TR=s ST/sinR = TR/sinS (4y + 6/sin 73 ) = (7y - 21/sin 73) so 4y + 6 = 7y-21 so 3y =27 y=9

OpenStudy (anonymous):

use sine rule here sin S / TR = sin R / ST substitute for these and solve for y